# Small Language Models: Survey, Measurements, and Insights

### v1

@@ 5 Conclusions and Future Directions
Device-cloud SLM-LLM collaboration. While our measurements demonstrate that SLM capability is fast evolving, the gap between it and the cloud-side LLM will exist. To achieve full-scale intelligence while not comprising privacy and availability (much), device-cloud collaboration will become an important research topic [17]. Intuitively, the SLM can be used as a “filter” that solves the easy tasks confidently on devices, and cloud LLM can be treated as a safe guard for critical, difficult tasks. This approach bottlenecks at the capability of SLM and the decision module of what tasks are easy enough for SLMs. Does a better collaboration approach exist? Such research is challenged by the auto-regressive manner of casual language models.

### v2

@@ 5 Conclusions and Future Directions
Device-cloud SLM-LLM collaboration. While our measurements demonstrate that SLM capability is fast evolving, the gap between it and the cloud-side LLM will exist. To achieve full-scale intelligence while not comprising privacy and availability (much), device-cloud collaboration will become an important research topic [19, 104]. Intuitively, the SLM can be used as a “filter” that solves the easy tasks confidently on devices, and cloud LLM can be treated as a safe guard for critical, difficult tasks. This approach bottlenecks at the capability of SLM and the decision module of what tasks are easy enough for SLMs. Does a better collaboration approach exist? Such research is challenged by the auto-regressive manner of casual language models.